HNMR for the given molecule. Peaks for specific functional groups.
200MH Ht 10 in eddod (ret, to CD00 e 330 pem OH Ho 7.0 6.9 6.8 Ppm 5 4 3 ppm 7 1.98 0.96 1.94 3.00
Answer #1

The peak of 1H of phenol is not shown here due to ionization.

g. @s Peaks s=3.8, singlet, 3H - s = 4.5, singlet, an - s = 6.71, doublet, Il - S= 6-76, doublet, 1 - S = 6.94, Singlet, 14